OpenStudy (anonymous):

What is the interval notation? x+1/x^2+2x-35

OpenStudy (anonymous):

\[\frac{ x+1 }{ x^2+2x-35 }=\frac{ x+1 }{ x^2+7x-5x-35 }=\frac{ x+1 }{ x \left( x+7\right)-5\left( x+7 \right) }\] \[=\frac{ x+1 }{\left( x+7 \right)\left( x-5 \right) }\] it is defined at all points except at x=-7,5 domain is \[x<-7,-7<x<5,x>5~or~x \in \left(- \infty,-7 \right)\cup \left( -7,-5 \right)\cup \left( 5,\infty \right)\]

OpenStudy (gorv):

second from last will be (-7,5)